OpenStudy (anonymous):

Suppose you need to solve a system of two equations in which both equations represent circles. How many solutions could your system have?

OpenStudy (anonymous):



OpenStudy (anonymous):

 technically you could have infinite if both equations are the same.

OpenStudy (anonymous):

I thought you were asking for example, if i have one circle with the formula y^2 + x^2 = c , and another with the formula y^2 + y^2 = d, where c and d are constants, how many solutions are there

OpenStudy (anonymous):

y^2 + x^2 = d for formula 2 sorry